# OmniPred: Language Models as Universal Regressors

### 23 Feb 2024 v2

@@ 2 Related Work and Motivation
Tensor representations are inherently constraint-dependent, as each tensor element must be in a reasonable numerical range (e.g. in $[-1,1]$) as inputs to a model. Thus to represent $x$, every scalar feature must be normalized against user-provided bounds, and every categorical feature must be one-hot embedded against user-provided choices. Dynamic yet minor input space changes such as new bounds or additional categories, are incompatible with this static representation. To represent $y$, a raw objective in $\mathbb{R}$ must also be normalized, which can be problematic at test-time when encountering outlier $y$-values. Dealing with this issue leads to implementing complicated nonlinear warpings (Daimon, 2011; Yeo & Johnson, 2000), many of which are also data-dependent (e.g. require storing min/max values from training data).
Regressor Dynamic Input Spaces? Can Multitask? Normalize $x$? Normalize $y$?
@@
Tree-based No Only fixed spaces Yes No
@@ 5.1 Simultaneous Regression
Figure 3: Model prediction samples over selected 4D shifted BBOB functions. Empirical mode (bolded) and min/max are shown from 10 samples. Over all BBOB functions, we vary the coordinate value $x_{i}$ while keeping others $x_{j\neq i}$ fixed.
@@
Figure 4: Left: Diagonal fit (/) is better. Model’s prediction vs. ground truth target objective over varying studies. Corporate-specific objective names are redacted. Right: Corresponding input spaces. “#-H, \$-T” is shorthand for a conditional input space with # root parameters and \$ total possible parameters.
@@ 5.2 Multi-task Transferrability
In Figure 6, we further see that for the model, multi-task training consistently improves over single-task training, and in regimes with relatively lower input space saturation from training data, multi-task models outperform traditional baselines over several different domains. Interestingly, a single-task model trained from scratch remains a competitive choice and for certain domains, can even outperform all other single-task baselines.
@@ 6 Conclusion
7 Limitations and Future Work
@@ B.1 Pretraining
The model ($\approx$ 200M parameters) was pretrained using a 4x4 TPU V2.
@@ B.2 Local Training
Due to the small training set and relatively low finetuning steps, we used a single 1x1 TPU V2.
@@ B.3 Inference
At inference time, we perform temperature sampling with a temperature of 1.0. We restrict the logits to only decode the custom floating point tokens for representing $y$-values. To maximize batch size for a 1x1 TPU V2, we generate 64 samples and select the empirical median of these floating point samples as our final prediction when computing error.
@@ C.4 Regressor Baselines
Tree and Random Forest: We use the standard API (XGBRegressor, XGBRFRegressor) found in XGBoost (https://github.com/dmlc/xgboost).

### 04 Mar 2024 v3

@@ 2 Related Work and Motivation
Tensor representations are commonly constraint-dependent, where each tensor element may need to be in a reasonable numerical range (e.g. in $[-1,1]$) as inputs to a model. Thus to represent $x$, every categorical feature must be one-hot embedded against user-provided choices, and scalar features may need to be rescaled against user-provided bounds. Dynamic yet minor input space changes such as new bounds or additional categories, are incompatible with this static representation. To represent $y$, a raw objective in $\mathbb{R}$ may also need to be rescaled, which can be problematic at test-time when encountering outlier $y$-values. Dealing with this issue leads to implementing complicated nonlinear warpings (Daimon, 2011; Yeo & Johnson, 2000), many of which are also data-dependent (e.g. require storing min/max values from training data).
Regressor Dynamic Input Spaces? Can Multitask? Tensorize? Rescale?
@@
Tree-based No Only fixed spaces Yes Optional
@@ 5.1 Simultaneous Regression
Figure 3: Model prediction samples over selected 4D BBOB functions with unseen shifts. Empirical mode (bolded) and min/max are shown from 10 samples. Over all BBOB functions, we vary the coordinate value $x_{i}$ while keeping others $x_{j\neq i}$ fixed.
@@
Figure 4: Left: Diagonal fit (/) is better. Model’s prediction vs. ground truth objective over varying studies. Corporate-specific objective names are redacted. Right: Corresponding input spaces. “#-H, \$-T” is shorthand for a conditional hybrid input space with # root parameters and \$ total possible parameters.
@@ 5.2 Multi-task Transferrability
In Figure 6, we further see that for the model, multi-task training consistently improves over single-task training, and in regimes with relatively lower input space saturation from training data, multi-task models outperform traditional baselines over several different domains. Interestingly, a single-task model trained from scratch remains a competitive choice and for certain domains such as AutoML, can even outperform all other single-task baselines. We hypothesize this is due to language-based representations being more appropriate for the conditional structures of these domains (e.g. AutoML).
@@ 6 Conclusion
7 Limitations and Possible Extensions
In this work, our emphasis was to demonstrate the promise of applying language modelling to general-purpose regression, and thus our design choices remained relatively simple to avoid confounding factors. We list some limitations of our specific design, which opens many more potential areas of exploration.
@@ B.1 Pretraining
The model ($\approx$ 200M parameters) was pretrained using a 4x4 TPU V3.
@@ B.2 Local Training
Due to the small training set and relatively low finetuning steps, we used a single 1x1 TPU V3.
@@ B.3 Inference
At inference time, we perform temperature sampling with a temperature of 1.0. We restrict the logits to only decode the custom floating point tokens for representing $y$-values. To maximize batch size for a 1x1 TPU V3, we generate 64 samples and select the empirical median of these floating point samples as our final prediction when computing error.
@@ C.4 Regressor Baselines
Below, we list out the specific implementation details of our regressor baselines. One nuanced issue is of hyperparameter tuning the regressors themselves, which could affect results. In order to be reasonably fair to all regressors (including our own OmniPred which has its own hyperparameters), for each regressor, we used a reasonable fixed set of hyperparameters for consistent throughout all experiments.
We emphasize in our paper on the concept of flexible string-based regression using language models and large-scale multi-task training data, rather than claiming replacement over widely accepted baselines in single-task, apples-to-apples comparisons.
@@
Tree and Random Forest: We use the standard API (XGBRegressor, XGBRFRegressor) found in XGBoost (https://github.com/dmlc/xgboost) with default hyperparameters. Although tree-based methods do not generally require rescaling, we still applied consistent $x$-preprocessing (in particular to deal with optional log or reverse-log scaling).